OpenStudy (anonymous):

Reduce to lowest terms. n^2 - 7n -44 _____________ n^2 -121 a) n -4/ n + 11 b) n + 11/ n + 4 c) n + 4/ n + 11

OpenStudy (anonymous):

n+4/n+11

OpenStudy (anonymous):

(n^2-11n+4n-44)/(n^2-121) =n(n-11)+4(n-11)/(n+11)(n-11) =(n+4)(n-11)/(n+11)(n-11) =n+4/n+11